# Breaking the Bulkhead: Demystifying Cross-Namespace Reference Vulnerabilities in Kubernetes Operators

### v1

@@ Abstract.
To the best of our knowledge, this paper is the first to systematically investigate the security vulnerability of Kubernetes Operators. We present Cross-Namespace Reference Vulnerability with two strategies, demonstrating how an attacker can bypass namespace isolation. Through large-scale measurements, we found that over 14% of Operators in the wild are potentially vulnerable. Our findings have been reported to the relevant developers, resulting in 7 confirmations and 6 CVEs by the time of submission, affecting vendors including the ****** and ******, highlighting the critical need for enhanced security practices in Kubernetes Operators. To mitigate it, we also open-source the static analysis suite to benefit the ecosystem.
Kubernetes, Operator, Vulnerability, Privilege Escalation, Cross-Namespace
copyright: noneccs: Security and privacy Software and application securityccs: Security and privacy Distributed systems security
1. Introduction
@@
Kubernetes has emerged as the dominant platform for container orchestration, playing a central role in the deployment, scaling, and management of containerized applications in modern cloud-native environments (cnc, 2024; red, 2024; edg, 2024; Team, 2024). As a highly extensible and open-source system, Kubernetes facilitates the automation of complex operations such as container deployment, scheduling, and management across clusters. Its flexibility and wide adoption have made it the cornerstone of many enterprise-level infrastructure solutions, offering efficient ways to handle diverse and dynamic workloads in a scalable manner.
@@
Kubernetes organizes resources into namespaces (nam, 2024), which allow users to divide a single cluster into multiple virtual clusters. Each namespace serves as a logical boundary, isolating resources like pods, services, and secrets from other namespaces within the same cluster. This isolation is essential for managing different applications or services within the same Kubernetes environment, enabling teams to work independently without interfering with each other. Namespaces also provide a way to scope access to resources, ensuring that certain actions can be confined to specific namespaces and reducing the risk of accidental or malicious interference between services.
@@
To achieve namespace isolation, a crucial security mechanism is Role-Based Access Control (RBAC) (rba, 2025). RBAC defines roles and permissions for users, service accounts, and other entities within the cluster, helping to control which actions are allowed within the system. For example, a staff member of a team may only be allowed to manipulate resources within the namespace assigned to their team, while the cluster administrator would be assigned all permissions across namespaces. By configuring RBAC policies, administrators can limit access to resources within specific namespaces, ensuring that users or services can only interact with the resources they are authorized to access. This granularity of access control reinforces the isolation between namespaces and helps to prevent unauthorized access to sensitive resources.
@@
While Kubernetes provides robust tools and security mechanisms to manage and secure applications, the native platform has limitations in automating the lifecycle of complex applications. Kubernetes requires significant manual intervention for tasks like scaling, upgrades, and configuration management, which can be time-consuming and error-prone (Vmware, 2022). Kubernetes Operators (ope, 2024b) were introduced to address these limitations. Operators are programs that extend Kubernetes’ capabilities by automating the management of applications. They encapsulate the operational knowledge required to manage complex Kubernetes applications, automating critical tasks such as deployment, scaling, and lifecycle management. Users can then easily request Operators to conduct complex operational tasks in contrast to manually manipulating raw Kubernetes resources. By automating these processes, Operators reduce the operational burden on DevOps teams and enable more consistent and reliable application management.
@@
Although extensive research has been conducted on Kubernetes security, the issue of Operator security remains largely unexplored. Previous studies have focused on misconfiguration in Kubernetes, especially excessive RBAC permissions, highlighting the risks of overly permissive access controls (Yang et al., 2023; Gu et al., 2025). While these works have led to improvements in reducing permissions, they have not addressed the inherent risks that remain even when permissions are minimized. Specifically, necessary permissions may still be exploited due to improper security practices within the Operator itself. Other research has focused on bugs in Kubernetes Operators (Xu et al., 2024; Sun et al., 2022; Gu et al., 2023), yet these studies primarily concentrated on the functional bugs of Operators rather than their security vulnerabilities. Furthermore, existing security tools (tri, 2025; kub, 2025d, a, b; opa, 2025; kyv, 2025) do not adequately address the security concerns specific to Operators, leaving a significant gap in the ecosystem.
Thus, in this paper, we present the first in-depth research on the security vulnerabilities of Kubernetes Operators, focusing on the risks of cross-namespace reference vulnerability. The root cause of cross-namespace reference vulnerability lies in the mismatch between the declared scope of a resource and the effective scope of the Operator’s logic. A resource may be declared as namespace-scoped, allowing users with limited access to a single namespace to deploy it, while the Operator’s logic may perform actions that affect other namespaces, breaking the intended isolation. We introduce two distinct tactics for exploiting Kubernetes Operators to elevate an attacker’s privileges, both of which exploit the scope mismatch in Operator implementation and break the isolation between Kubernetes namespaces.
To aid in detecting these vulnerabilities, we designed and implemented detection tools that can identify scope mismatch in Operators, enabling users to assess and secure their systems against cross-namespace attacks. We conducted large-scale measurements of 2,268 Kubernetes Operators in the wild, revealing that over 14% of the Operators are potentially vulnerable to these attacks. We responsibly disclosed our findings to their developers, and, by the time of submission, 7 vulnerabilities had been confirmed and 6 CVEs were assigned or under assignment in response to our reports, affecting vendors including ****** and ******, highlighting the critical need for enhanced security practices in Kubernetes Operators.
@@
  • Attack Surface. We present the first systematic research on Kubernetes Operator security, specifically focusing on vulnerabilities that can be exploited in Operator implementations.
  • Attack Tactics. We introduce the Cross-Namespace Reference Vulnerability, detailing two distinct tactics and their root cause, both of which can be leveraged to elevate an attacker’s privileges.
  • Automatic Detection. We design and implement tools for detecting vulnerabilities in Kubernetes Operators, enabling users to identify risks related to cross-namespace reference vulnerability.
  • Real-World Measurement. We conduct large-scale measurements of Kubernetes Operators in the wild, demonstrating that over 14% of Operators are susceptible to these vulnerabilities.
  • Responsible Disclosure. We responsibly disclosed our findings to the developers, and, by the time of submission, 7 vulnerabilities were confirmed and 6 CVEs were assigned or under assignment in response to our reports.
  • Benefit Ecosystem. We open-source our tools to the community at https://anonymous.4open.science/r/Operator-Vuln-F05C, thereby fostering greater security awareness and improving the overall Kubernetes ecosystem.
2. Background
2.1. Kubernetes Namespace and RBAC
Kubernetes is a powerful container orchestration platform that automates the deployment, scaling, and management of containerized applications. It is designed to manage large-scale complex applications, where multiple teams or applications may share a single cluster. To help organize and isolate resources within the cluster, Kubernetes provides a mechanism called Namespaces (nam, 2024). A namespace is a logical partition or a virtual cluster within a physical cluster. Each namespace acts as a boundary, ensuring that resources in one namespace do not conflict with those in another.
Namespaces are particularly useful in multi-tenant environments, where different teams or applications share the same Kubernetes cluster (kub, 2024). By isolating resources in separate namespaces, Kubernetes prevents one team from accessing or interfering with another team’s resources. This isolation is vital for security and resource management, ensuring that users and applications can only access the resources assigned to their namespace, preventing unauthorized access or potential conflicts between resources.
Kubernetes employs multiple security mechanisms to help ensure that the cluster remains secure and that resources are properly isolated. One of the most important mechanisms for securing access to resources within a namespace is RBAC (Role-Based Access Control) (rba, 2025). RBAC allows administrators to define roles with specific permissions and bind those roles to users or service accounts, ensuring that only authorized entities can perform certain actions. Cluster administrator can grant both namespace-specific permissions and cluster-level permissions. Specifically:
@@ 2.1. Kubernetes Namespace and RBAC
2.2. Kubernetes Resource
At its core, Kubernetes organizes the cluster’s state using resources, which are data objects encapsulating configuration and runtime information. Kubernetes manages many types of resources within a cluster, which are fundamental components that define the desired state of applications and services. Common built-in resources types include pods (pod, 2025) and deployments (dep, 2025). A pod is the smallest deployable unit in Kubernetes and typically represents one or more containers that share the same network and storage resources. A deployment is a higher-level abstraction that manages the lifecycle of pods, specifying the desired number of pod replicas. In addition to built-in resources, Kubernetes allows users to define Custom Resources (CRs) (cr, 2024), extending Kubernetes to manage domain-specific requirements beyond its default capabilities. Each type of Custom Resource is described by a Custom Resource Definition (CRD) (cr, 2024), which specifies the resource’s schema.
In Kubernetes, each type of resource, whether a built-in resource or a Custom Resource, is bound with an explicit scope, indicating its accessibility and impact within the cluster. Resources can be either Namespace-scoped or Cluster-scoped. In Kubernetes, the scope of built-in resources is embedded within the Kubernetes implementation, whereas the scope of Custom Resources is explicitly defined in their associated Custom Resource Definitions. Namespace-scoped resources must reside within a specific namespace, meaning they are logically isolated and can be accessed or manipulated by users with only Namespace-specific Roles. Conversely, Cluster-scoped resources exist at the cluster level and are not confined to any single namespace. These Cluster-scoped resources may affect or interact with all namespaces across the cluster. Due to their broad impact, accessing or manipulating cluster-scoped resources requires users to possess a cluster-wide ClusterRole, reflecting elevated privileges.
@@ 2.2. Kubernetes Resource
2.3. Kubernetes Operator
A Kubernetes Operator is a method of automating and managing the lifecycle of complex applications on top of Kubernetes by extending the platform’s native capabilities. Originally introduced by CoreOS (now part of Red Hat), it emerged from a recognition that while Kubernetes excels at automatically orchestrating workloads, many organizations need a more powerful automation pattern to handle full lifecycle management, such as database management, application upgrades, or failure recovery, that requires specific operational knowledge. Operator is thus introduced to extend Kubernetes by embedding human operational expertise into software, enabling automated management of complex, stateful applications.
An Operator consists of one or more Custom Resource Definitions (CRDs) and their corresponding Custom Resource Controllers. CRD defines the schema of a custom resource type that will be processed by the Operator. Controllers work with CRDs by continually monitoring the custom resources and taking actions to fulfill the operational tasks requested by users.
To use an Operator, users manipulate custom resources that represent the operational task along with the related parameters they want to conduct. The Operator controller reads these custom resources, takes actions listed in the custom resource, and ensures that the operational task behaves as expected.
Figure 1. Custom Resource Example
[…]
Custom Resource Example
@@ 2.3. Kubernetes Operator
3. Threat Model
Our threat model aligns with real-world Kubernetes deployments where multiple tenants, teams, or applications share the same cluster while being isolated within their respective namespaces (kub, 2024). The adversary aims to break Kubernetes namespace isolation and achieve cross-namespace privilege escalation by exploiting security weaknesses in Operator implementations. Their objectives are performing operations in unauthorized namespaces (i.e., namespaces that they have no Roles) and thus escalating privileges.
We assume the Kubernetes cluster deploys vulnerable Operators, and the adversary has legitimate access to a Kubernetes cluster but can only access their authorized namespaces. Thus, they can not access or manipulate cluster-scoped resources and can only interact with Operators by manipulating namespace-scoped resources in their authorized namespace. They seek to leverage vulnerable Kubernetes Operators to execute unauthorized operations in other namespaces. The adversary may be:
@@ 3. Threat Model
  • A compromised application running in a namespace with namespace-level permissions mounted.
It is notable that our threat model is significantly different from previous works (Yang et al., 2023; Gu et al., 2025). Specifically, existing works assume that the adversaries have direct control over the vulnerable application containers, which is a strong assumption in the real-world. In contrast, within our threat model, an adversary does not have access to Kubernetes work nodes or direct control of Operator containers. They cannot directly access containers of Operators since these Operators may be deployed in adversary-unauthorized namespaces. In extreme cases, Operators can even be deployed outside the Kubernetes cluster (ope, 2024a). So the threat model of previous works (Yang et al., 2023; Gu et al., 2025) is relatively infeasible, but our threat model is more feasible and aligned with real-world scenarios.
4. Cross-Namespace Attacks
Figure 2. Attack Flow
4.1. Attack Overview
@@
The high-level attack flow is as follows: an attacker, who has legitimate but restricted access to one namespace, manipulates a maliciously crafted namespace-scoped resource instance within their authorized namespace. The Operator, continuously watching for namespace-scoped resource events, detects this malicious resource instance and processes it with privileged operations that impact namespaces beyond the attacker’s authorized scope, effectively breaking the intended namespace isolation enforced by Kubernetes.
Root Cause. The core enabling cross-namespace reference attacks stems from a mismatch between the declared scope of a resource and the actual scope of its process logic. Specifically, the vulnerability arises when the scope of a resource is defined as Namespaced, indicating that each instance should strictly reside within its assigned namespace. Thus an adversary only with Role in a single namespace may manipulate such a resource in their namespace. However, despite this namespace-scoped definition, the Operator may actually perform operations across namespaces, inadvertently allowing manipulation of resources in namespaces beyond the intended scope. As a result, an adversary without Role in other namespaces may invoke the Operator to escalate their permissions and access unauthorized namespaces.
Cross-Namespace Features. There are two primary scenarios that enable cross-namespace reference actions. First, when processing namespace-scoped resources in one namespace, an Operator may access or manipulate other namespace-scoped resources in a different namespace (§4.2). Second, when processing namespace-scoped resources, an Operator might access or manipulate cluster-scoped resources, leading to impacts on the whole cluster across all namespaces (§4.3). Both scenarios allow adversaries to trick the Operator into performing unintended, privileged operations beyond the adversary’s RBAC scope.
4.2. Insecure Namespace-Scoped Resource Reference
Insecure Namespace-Scoped Resource Reference vulnerability arises when an Operator processing namespace-scoped resources, the fields of which are then used by the Operator to reference resources in other namespaces. This vulnerability fundamentally undermines Kubernetes namespace isolation by enabling attackers to indirectly access resources from namespaces they are otherwise restricted from accessing.

### v2

@@ Abstract
To the best of our knowledge, this paper is the first to systematically investigate Kubernetes Operator attacks. We present Cross-Namespace Reference Vulnerability with two strategies, demonstrating how an attacker can bypass namespace isolation. Through large-scale measurements, we found that over 14% of Operators in the wild are potentially vulnerable. Our findings have been reported to the relevant developers, resulting in 8 confirmations and 7 CVEs by the time of submission, affecting vendors including Red Hat and NVIDIA, highlighting the critical need for enhanced security practices in Kubernetes Operators. To mitigate it, we open-source the static analysis suite and propose concrete mitigation to benefit the ecosystem.
publicationid: pubid: ❡Corresponding Authors: Ziyi Guo, Zhenyuan Li Network and Distributed System Security (NDSS) Symposium 2026 23 - 27 February 2026, San Diego, CA, USA ISBN 979-8-9919276-8-0 https://dx.doi.org/10.14722/ndss.2026.240761 www.ndss-symposium.org
I Introduction
@@
Kubernetes has emerged as the dominant platform for container orchestration, playing a central role in the deployment, scaling, and management of containerized applications in modern cloud-native environments [cncfreport, redhatreport, edgedelta, datahub]. As a highly extensible and open-source system, Kubernetes facilitates the automation of complex operations such as container deployment, scheduling, and management across clusters. Its flexibility and wide adoption have made it the cornerstone of many enterprise-level infrastructure solutions, offering efficient ways to handle diverse and dynamic workloads in a scalable manner.
@@
Kubernetes organizes resources into namespaces [namespace], which allow users to divide a single cluster into multiple virtual clusters. Each namespace serves as a logical boundary, isolating resources like pods, services, and secrets from other namespaces within the same cluster. This isolation is essential for managing different applications or services within the same Kubernetes environment, enabling users to work independently without interfering with each other. Namespaces also provide a way to scope access to resources, ensuring that certain actions can be confined to specific namespaces and reducing the risk of accidental or malicious interference between services.
@@
To achieve namespace isolation, a crucial security mechanism is Role-Based Access Control (RBAC) [rbac]. RBAC defines roles and permissions for users, service accounts, and other entities within the cluster, helping to control which actions are allowed within the system. For example, a staff member of a team may only be allowed to manipulate resources within the namespace assigned to their team, while the cluster administrator would be assigned all permissions across namespaces. By configuring RBAC policies, administrators can limit access to resources within specific namespaces, ensuring that users or services can only interact with the resources they are authorized to access. This granularity of access control reinforces the isolation between namespaces and helps to prevent unauthorized access to sensitive resources.
@@
While Kubernetes provides robust tools and security mechanisms to manage and secure applications, the native platform has limitations in automating the lifecycle of complex applications. Kubernetes requires significant manual intervention for tasks like scaling, upgrades, and configuration management, which can be time-consuming and error-prone [vmwarereport]. Kubernetes Operators [operator] were introduced to address these limitations. Operators are programs that extend Kubernetes’ capabilities by automating the management of applications. They encapsulate the operational knowledge required to manage complex Kubernetes applications, automating critical tasks such as deployment, scaling, and lifecycle management. Users can then easily request Operators to conduct complex operational tasks in contrast to manually manipulating raw Kubernetes resources. By automating these processes, Operators reduce the operational burden on DevOps teams and enable more consistent and reliable application management.
@@
Although extensive research has been conducted on Kubernetes security, Operator-specific attacks remain largely unexplored. Previous studies have focused on misconfiguration in Kubernetes, especially excessive RBAC permissions, highlighting the risks of overly permissive access controls [takeover, epscan, podescape, k8sblackhat]. Practitioners thus suggested limiting RBAC permissions for Kubernetes Operators [operatorsecpractice, operatorsynk, operatorkubeops]. While these works have led to improvements in reducing permissions, they have neither addressed the inherent risks that remain even when permissions are minimized, nor presented attacks specific to Operators. Specifically, necessary permissions required by Operator business logics, which cannot be further minimized, may still be exploited due to improper security design within the Operator logic. Furthermore, existing attacks assume attackers have compromised containers (e.g., get a shell) in prior, leaving critical gaps in how to compromise application containers.
Other research has focused on bugs in Kubernetes Operators [operatorbug, sieve, acto], yet these studies primarily concentrated on the functional bugs of Operators rather than their security vulnerabilities. Furthermore, existing security tools [trivy, kubescape, kubesec, kubearmor, opa, kyverno] do not adequately address the security concerns specific to Operators, leaving a significant gap in the ecosystem.
Thus, in this paper, we present the first systematic research on Kubernetes Operator attacks, unveiling cross-namespace reference vulnerabilities. The root cause of cross-namespace reference vulnerability lies in the mismatch between the declared scope of a resource and the effective scope of the Operator’s logic. A resource may be declared as namespace-scoped, allowing users with limited access to a single namespace to deploy it, while the Operator’s logic may perform actions that affect other namespaces, breaking the intended isolation. We present a novel practical threat model without assuming that attackers have already compromised containers, allowing exploits to be made from scratch. We propose two distinct tactics for exploiting Kubernetes Operators to elevate an attacker’s privileges, both of which exploit the scope mismatch in Operator implementation and break the isolation between Kubernetes namespaces.
To measure the new kind of vulnerabilities, we designed and implemented a static analysis suite that can identify scope mismatch in Operators. We conducted large-scale measurements of 2,268 Kubernetes Operators in the wild, revealing that over 14% of the Operators are potentially vulnerable to these attacks. We responsibly disclosed our findings to their developers, and, by the time of submission, 8 vulnerabilities had been confirmed and 7 CVEs were assigned or under assignment in response to our reports, affecting vendors Red Hat and NVIDIA, highlighting the critical need for enhanced security practices in Kubernetes Operators.
@@
  • New Attack. We present the first systematic research on Kubernetes Operator attacks. We unveil the new type of vulnerability specific to Operators, Cross-Namespace Reference Vulnerability, detailing two distinct tactics and their root cause, both of which can be leveraged to escalate privileges.
  • Large-scale Measurement. We design and implement tools for the measurement of real-world Cross-Namespace Reference Vulnerabilities. We conduct large-scale measurements of Kubernetes Operators in the wild, demonstrating that over 14% of Operators are susceptible to these vulnerabilities.
  • Real-World Impact. We responsibly disclosed our findings to the developers, and, by the time of submission, 8 vulnerabilities were confirmed and 7 CVEs were assigned in response to our reports, affecting leading vendors including Red Hat and NVIDIA.
  • Benefic Ecosystem. We will open-source our analyzer, which covers detailed modeling of major Kubernetes libraries, enabling analysis of not only Operators but also other Kubernetes applications. To facilitate remediation, we propose concrete mitigations, open-source mitigation samples, and a patch generator for practitioners.
II Background
II-A Kubernetes Namespace and RBAC
Kubernetes is a powerful container orchestration platform that automates the deployment, scaling, and management of containerized applications. It is designed to manage large-scale complex applications, where multiple teams or applications may share a single cluster. To help organize and isolate resources within the cluster, Kubernetes provides a mechanism called Namespaces [namespace]. A namespace [^1] We use ns as short for namespace for the remaining parts. is a logical partition or a virtual cluster within a physical cluster. Each namespace acts as a boundary, ensuring that resources in one namespace do not conflict with those in another.
Namespaces are particularly useful in multi-tenant environments, where different teams or applications share the same Kubernetes cluster [kubernetes-multitenancy]. By isolating resources in separate namespaces, Kubernetes prevents one team from accessing or interfering with another team’s resources. This isolation is vital for security and resource management, ensuring that users and applications can only access the resources assigned to their namespace, preventing unauthorized access or potential conflicts between resources.
Kubernetes employs multiple security mechanisms to help ensure that the cluster remains secure and that resources are properly isolated. One of the most important mechanisms for securing access to resources within a namespace is RBAC (Role-Based Access Control) [rbac]. RBAC allows administrators to define roles with specific permissions and bind those roles to users or service accounts, ensuring that only authorized entities can perform certain actions. Cluster administrators can grant both ns-specific permissions and cluster-level permissions. Specifically:
@@ II-A Kubernetes Namespace and RBAC
II-B Kubernetes Resource
At its core, Kubernetes organizes the cluster’s state using resources, which are data objects encapsulating configuration and runtime information. Kubernetes manages many types of resources within a cluster, which are fundamental components that define the desired state of applications and services. Common built-in resource types include pods [pod] and deployments [deployment]. A pod is the smallest deployable unit in Kubernetes and typically represents one or more containers that share the same network and storage resources. A deployment is a higher-level abstraction that manages the lifecycle of pods, specifying the desired number of pod replicas. In addition to built-in resources, Kubernetes allows users to define Custom Resources (CRs) [cr], extending Kubernetes to manage domain-specific requirements beyond its default capabilities. Each type of Custom Resource is described by a Custom Resource Definition (CRD) [cr], which specifies the resource’s schema.
In Kubernetes, each type of resource, whether a built-in resource or a Custom Resource, is bound with an explicit scope, indicating its accessibility and impact within the cluster. Resources can be either Namespace-scoped or Cluster-scoped. In Kubernetes, the scope of built-in resources is embedded within the Kubernetes implementation, whereas the scope of Custom Resources is explicitly defined in their associated Custom Resource Definitions. NS-scoped resources must reside within a specific namespace, meaning they are logically isolated and can be accessed or manipulated by users with only NS-specific Roles. Conversely, Cluster-scoped resources exist at the cluster level and are not confined to any single namespace. These Cluster-scoped resources may affect or interact with all namespaces across the cluster. Due to their broad impact, accessing or manipulating cluster-scoped resources requires users to possess a cluster-wide ClusterRole, reflecting elevated privileges.
@@ II-B Kubernetes Resource
II-C Kubernetes Operator
Kubernetes Operator is a method of automating and managing the lifecycle of complex applications on top of Kubernetes by extending the platform’s native capabilities. Originally introduced by CoreOS (now part of Red Hat), it emerged from a recognition that while Kubernetes excels at automatically orchestrating workloads, many organizations need a more powerful automation pattern to handle full lifecycle management, such as database management, application upgrades, or failure recovery, that requires specific operational knowledge. Operator is thus introduced to extend Kubernetes by embedding human operational expertise into software, enabling automated management of complex, stateful applications.
An Operator consists of one or more Custom Resource Definitions (CRDs) and their corresponding Custom Resource Controllers. CRD defines the schema of a custom resource type that will be processed by the Operator. Controllers work with CRDs by continually monitoring custom resources and taking actions to fulfill operational tasks requested by users.
To use an Operator, users manipulate custom resources that represent the operational task, along with the related parameters they want to conduct. The Operator controller reads these custom resources, takes actions listed in the custom resource, and ensures that the operational task behaves as expected.
Figure 1: Custom Resource Example
@@ II-C Kubernetes Operator
III Threat Model
Our threat model aligns with real-world Kubernetes deployments where multiple tenants, teams, or applications share the same cluster while being isolated within their respective namespaces [kubernetes-multitenancy]. The adversary aims to break Kubernetes namespace isolation and achieve cross-ns privilege escalation by exploiting security weaknesses in Operator implementations. Their objectives are performing operations in unauthorized namespaces (i.e., namespaces that they have no Roles) and thus escalating privileges.
We assume the Kubernetes cluster deploys vulnerable Operators, and the adversary has legitimate access to a Kubernetes cluster but can only access their authorized namespaces. Thus, they cannot access or manipulate cluster-scoped resources and can only interact with Operators by manipulating ns-scoped resources in their authorized namespace. They seek to leverage vulnerable Kubernetes Operators to execute unauthorized operations in other namespaces. The adversary may be:
@@ III Threat Model
  • A compromised application running in a namespace with ns-level permissions mounted.
  • An attacker who steals credentials of Kubernetes accounts with ns-level permissions.
Our threat model is practical. For example, many real-world cloud services are operated on multi-tenant Kubernetes. One of them, Red Hat Developer Sandbox [developersandbox], is provided by assigning a namespace on multi-tenant Kubernetes to a user, where Operators are deployed. This implies that attackers may subscribe Kubernetes-based service to conduct our attacks.
It is notable that our threat model is significantly different from previous works [takeover, epscan, podescape, k8sblackhat]. Specifically, existing works assume that the adversaries have compromised vulnerable application containers in prior, which is a strong assumption in the real world, leaving critical gaps in how to compromise application containers. In contrast, within our threat model, adversaries don’t have to gain control of Operators in prior, since these Operators may be deployed in adversary-unauthorized namespaces. In extreme cases, Operators can even be deployed outside the Kubernetes cluster [operatoroutside]. So the threat model of previous works is relatively infeasible, but our threat model is more feasible and aligned with real-world scenarios.
In this paper, the terms Namespace and Cross-Namespace refer specifically to Kubernetes Namespaces, which are used to isolate resources within a Kubernetes cluster. They are distinct from similarly named concepts in other systems, such as Linux Namespaces, which isolate system resources at the OS kernel level. They isolate resources at different layers and have no direct linkage despite sharing the same term.
The vulnerability we presented acts as a strategy for attackers to access or manipulate unauthorized resources. The final concrete impact depends on the accessed unauthorized resources and differs between cases.
IV Cross-Namespace Attacks
Figure 2: Attack Flow
IV-A Attack Overview
@@
The high-level attack flow is as follows: an attacker, who has legitimate but restricted access to one namespace, manipulates a maliciously crafted ns-scoped resource instance within their authorized namespace. The Operator, continuously watching for ns-scoped resource events, detects this malicious resource instance and processes it with privileged operations that impact namespaces beyond the attacker’s authorized scope, effectively breaking the intended namespace isolation enforced by Kubernetes.
Root Cause. The core enabling cross-ns reference attacks stems from a mismatch between the declared scope of a resource and the actual scope of its process logic. Specifically, the vulnerability arises when the scope of a resource is defined as Namespaced, indicating that each instance should strictly reside within its assigned namespace. Thus, an adversary only with Role in a single namespace is allowed by RBAC to manipulate such a resource in their own namespace. However, despite this ns-scoped definition, the Operator may actually perform operations across namespaces, inadvertently allowing manipulation of resources in namespaces beyond the intended scope. As a result, an adversary without Role in other namespaces may invoke the Operator to escalate their permissions and access unauthorized namespaces.
The root leading to such implementation lies in the improper trade-off at the design level between security and convenience. Kubernetes grants significant flexibility for controllers. Built-in controllers in Kubernetes prioritize security to eliminate attack surfaces. For instance, a Pod can only reference Secrets within the same namespace to avoid cross-ns. However, many developers [kubeissue57325, kubeissue104415, kubeissue107495, wantcrossns, wantcrossns2] want cross-ns to avoid manually duplicating Secrets in each namespace. Since Operators are proposed to reduce human intervention, the community may prioritize convenience, overlooking the underlying Kubernetes’s security model and enabling cross-ns reference.
Cross-Namespace Features. There are two primary scenarios that enable cross-namespace reference actions. First, when processing ns-scoped resources in one namespace, an Operator may access or manipulate other ns-scoped resources in a different namespace (§IV-B). Second, when processing ns-scoped resources, an Operator might access or manipulate cluster-scoped resources, leading to impacts on the whole cluster across all namespaces (§IV-C). Both scenarios allow adversaries to trick the Operator into performing unintended, privileged operations beyond the adversary’s RBAC scope.
IV-B Insecure Namespace-Scoped Resource Reference
Insecure NS-Scoped Resource Reference vulnerability arises when an Operator processing ns-scoped resources, the fields of which are then used by the Operator to reference resources in other namespaces. This vulnerability fundamentally undermines Kubernetes namespace isolation by enabling attackers to indirectly access resources from namespaces they are otherwise restricted from accessing.